OpenStudy (anonymous):

nput the equation of the given line in standard form. The line including (3, 1) and (-2, 3).

sam (.sam.):

First find the slope

OpenStudy (anonymous):

I don't even know how to find the slope.

sam (.sam.):

\[\Huge \color{blue}{Slope}=\color{green}{\frac{y-y_1}{x-x_1}}\]

sam (.sam.):

\[\Huge \color{blue}{Slope}=\color{green}{\frac{1-3}{3-(-2)}}\] \[\Huge \color{blue}{Slope}=\color{green}{\frac{2}{5}}\]

sam (.sam.):

Then use \[\huge \mathbb{\color{green}{y-y_1=m(x-x_1)}}\]

sam (.sam.):

\[\huge \mathbb{\color{purple}{y-1=\frac{2}{5}(x-3)}}\] expand it

sam (.sam.):

\[-2x+5y+1=0\]